OpenStudy (anonymous):

2,6,18,54.....t10 ???

OpenStudy (raden):

this is the sequence of geometeri with the 1st term t1= 2 and its ratio r = t2/t1 = 6/2 = 3 to get the value of t10, use this formula : tn = t1*r^(n-1) so, t10 = 2 * 3^(10-1) = 2 * 3^9 = 2 * 19,683 = 39,366